# Regret Analysis of Posterior Sampling-Based Expected Improvement for Bayesian Optimization

### v2

@@ Appendix D Auxiliary Lemmas
Let $f\sim{\cal G}{\cal P}(0,k)$, where $k$ is a kernel s.t. $k(\boldsymbol{x},\boldsymbol{x})\leq 1$. If ${\cal X}$ is finite, $\mathbb{E}[\max_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})]=\mathbb{E}[\min_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})]\leq\sqrt{2\log\left(|{\cal X}|/2\right)+2}$. If Assumption 2.1 holds, $\mathbb{E}[\max_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})]=\mathbb{E}[\min_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})]\leq\sqrt{2d\log\left(\left\lceil bdr(\log(ad)+\sqrt{\pi}/2)\right\rceil\right)+2-2\log 2}+1$.
@@
First, since $f$ follows a centered GP, $\mathbb{E}[\max_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})]=\mathbb{E}[\min_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})]$. For a finite input domain ${\cal X}$, from Lemma 4.2 of [Takeno et al., 2023], we can obtain
@@
For an infinite input domain, we consider a finite set ${\cal X}_{t}$ with each dimension evenly divided by $\tau_{t}=\left\lceil bdr(\log(ad)+\sqrt{\pi}/2)\right\rceil$, such that $|{\cal X}_{t}|=\tau_{t}^{d}$. In addition, we denote $[\boldsymbol{x}]_{t}=\operatorname*{arg\,min}_{\boldsymbol{x}^{\prime}\in{\cal X}_{t}}\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|_{1}$. Then, by Lemma H.2 of [Takeno et al., 2023], we see that
@@
Figure 3: Results of simple regret for synthetic functions generated from GP defined by the Matérn kernel with $\nu=5/2$. The top and bottom rows show the result on the length scale of the kernel $\ell=0.1$ and $\ell=0.2$, respectively. The left, centered, and right columns show the result on the noise standard deviation $\sigma=0.01,0.1$, and $1$, respectively.
Figure 4: Results of cumulative regret for synthetic functions generated from GP defined by the Matérn kernel with $\nu=5/2$. The top and bottom rows show the result on the length scale of the kernel $\ell=0.1$ and $\ell=0.2$, respectively. The left, centered, and right columns show the result on the noise standard deviation $\sigma=0.01,0.1$, and $1$, respectively.
@@ E.2 Benchmark Function Experiments
Figure 5 shows the results of the simple regret for several benchmark functions. We employed the SE kernel function with automatic relevance determination [Rasmussen and Williams, 2005]. We performed hyperparameter selection for the GP model every five iterations. In these experiments, we did not add actual noise. Since $\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\mu_{t}(\boldsymbol{x})$ is unstable due to the hyperparameter selection, we define the simple regret using $\hat{\boldsymbol{x}}_{t}=\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\{\mu_{t}(\boldsymbol{x})-2\sigma_{t}(\boldsymbol{x})\}$. We employed the heuristic hyperparameters for the AFs of GP-UCB, IRGP-UCB, and GP-EI-$\mu^{\rm max}$ as $\beta_{t}=0.2d\log(2t)$ as with [Kandasamy et al., 2016], $\nu_{t}=0.2d\log(2t)$, and $\zeta_{t}=\max\{0.2d\log(2t)-2,0\}+Z_{t}$, where $Z_{t}\sim{\rm Exp}(\lambda=1/2)$. Therefore, GP-UCB, IRGP-UCB, and GP-EI-$\mu^{\rm max}$ are no longer theoretically guaranteed and indicated by dashed lines. Other settings are the same as those of the synthetic function experiments.
Except for the Ackley and Shekel functions, almost all BO methods perform well. For the Ackley function, GP-UCB, IRGP-UCB, and GP-EI show superior performance. This is because these BO methods are more exploitative than other BO methods in this experiment, and exploitative AFs work well for the Ackley function. However, for the Shekel function, more explorative GP-PIMS and GP-EIMS show superior performance. Therefore, we believe that these differences in performance come from the exploitation-exploration tradeoff. On the other hand, regarding the BO methods with the theoretical guarantee without heuristic tunings, we consider that GP-PIMS and GP-EIMS are superior or comparable to GP-TS.
Figure 5: Results of simple regret for the benchmark functions. The BO methods with the theoretical guarantee, that is, GP-EIMS, GP-PIMS, and GP-TS, are indicated by solid lines.

### v3

@@ Appendix D Auxiliary Lemmas
Let $f\sim{\cal G}{\cal P}(0,k)$, where $k$ is a kernel s.t. $k(\boldsymbol{x},\boldsymbol{x})\leq 1$. If $\cal X$} is finite, $\mathbb{E}[\max_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})]=-\mathbb{E}[\min_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})]\leq\sqrt{2\log\left(|{\cal X}|/2\right)+2}$. If Assumption 2.1 holds, $\mathbb{E}[\max_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})]=-\mathbb{E}[\min_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})]\leq\sqrt{2d\log\left(\left\lceil bdr(\log(ad)+\sqrt{\pi}/2)\right\rceil\right)+2-2\log 2}+1$.
@@
First, since $f$ follows a centered GP, $\mathbb{E}[\max_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})]=-\mathbb{E}[\min_{\boldsymbol{x}\in{\cal X}}f(\boldsymbol{x})]$. For a finite input domain $\cal X$}, from Lemma 4.2 of [Takeno et al., 2023], we can obtain
@@
For an infinite input domain, we consider a finite set ${\cal X}_{t}$ with each dimension evenly divided by $\tau_{t}=\left\lceil bdr(\log(ad)+\sqrt{\pi}/2)\right\rceil$, such that $|{\cal X}_{t}|=\tau_{t}^{d}$. In addition, we denote $\boldsymbol{x}]_{t}=\operatorname*{arg\,min}_{\boldsymbol{x}^{\prime}\in{\cal X}_{t}}\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|_{1}$. Then, by Lemma H.2 of [Takeno et al., 2023], we see that
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 3: Results of simple regret for synthetic functions generated from a GP defined by the Matérn kernel with $\nu=5/2$. The top, middle, and bottom rows show the result for noise standard deviations $\sigma=0.01,0.1$, and $1$, respectively. The left and right columns show the result for length scales of the kernel $\ell=0.1$ and $\ell=0.2$, respectively. Daggers in the legend indicate that the BO method was performed with theoretical settings, for example, regarding the hyperparameters.
@@ E.2 Benchmark Function Experiments
Figure 5 shows the results of the simple regret for several benchmark functions. We employed the SE kernel function with automatic relevance determination [Rasmussen and Williams, 2005]. We performed hyperparameter selection for the GP model every five iterations. In these experiments, we did not add actual noise. Since $\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\mu_{t}(\boldsymbol{x})$ is unstable due to the hyperparameter selection, we define the simple regret using $\hat{\boldsymbol{x}}_{t}=\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\{\mu_{t}(\boldsymbol{x})-2\sigma_{t}(\boldsymbol{x})\}$. We employed the heuristic hyperparameters for the AFs of GP-UCB, IRGP-UCB, and GP-EI-$\mu^{\rm max}$ as $\beta_{t}=0.2d\log(2t)$ as with [Kandasamy et al., 2016], $\nu_{t}=0.2d\log(2t)$, and $\zeta_{t}=\max\{0.2d\log(2t)-2,0\}+Z_{t}$, where $Z_{t}\sim{\rm Exp}(\lambda=1/2)$. Therefore, GP-UCB, IRGP-UCB, and GP-EI-$\mu^{\rm max}$ are no longer theoretically guaranteed. Other settings are the same as those of the synthetic function experiments.
Except for the Ackley and Shekel functions, almost all BO methods perform well. For the Ackley function, GP-UCB, IRGP-UCB, and GP-EI show superior performance. This is because these BO methods are more exploitative than other BO methods in this experiment, and exploitative AFs work well for the Ackley function. However, for the Shekel function, more explorative GP-PIMS, GP-EIMS, and E3I show superior performance. Therefore, we believe that these differences in performance come from the exploitation-exploration tradeoff. On the other hand, regarding the BO methods with the theoretical guarantee without heuristic tunings, we consider that GP-PIMS and GP-EIMS are superior or comparable to GP-TS.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: Results of cumulative regret for synthetic functions generated from a GP defined by the Matérn kernel with $\nu=5/2$. The top, middle, and bottom rows show the result for noise standard deviations $\sigma=0.01,0.1$, and $1$, respectively. The left and right columns show the result for length scales of the kernel $\ell=0.1$ and $\ell=0.2$, respectively. Daggers in the legend indicate that the BO method was performed with theoretical settings, for example, regarding the hyperparameters.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 5: Results of simple regret for the benchmark functions. Daggers in the legend indicate that the BO method was performed with theoretical settings, for example, regarding the hyperparameters.